# Network Formation and Dynamics Among Multi-LLMs

### v2

@@ Acknowledgements.
M.P. is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
@@ Appendix A Regression Table
In Table S.1, we report the regression coefficient for the regression in the real-world network data. The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 2.3 (also shown in Figure 6) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
@@ Appendix C LLM Reasoning Summarization
In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions. After LLMs form networks, we ask them about their reasoning. We then use GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) to group their explanations into common themes, creating various categories. To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
@@ C.1. Micro-level Principles
In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 2.1.
In Figures 7(a) and 7(b), the reasons mean the following:
  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases. This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment. For Figure 7(b), we observe an opposite pattern in the LLM reasoning. A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
Finally, Figure 7(c) contains the reasons for Figure 3(a) based on having a common hobby, a common color, or a common location. We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
@@
Figure S.7. Distribution of reasoning across categories for the networks of Figures 1(a), 2(a) and 3(b). The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies). The error bars correspond to 95% confidence intervals.

### v3

@@ Acknowledgements.
The authors would like to thank Jon Kleinberg, Yanbang Wang, Yuanqi Du, the attendees of the Learning on Graphs NYC meetup, and the participants of Cornell AI, Policy, and Practice working group for their valuable feedback on the current version of the paper. M.P. is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
@@ Appendix A Regression Table
In Table S.1, we report the regression coefficient for the regression in the real-world network data. The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
@@ Appendix C LLM Reasoning Summarization
In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions. After LLMs form networks, we ask them about their reasoning. We then use GPT models (GPT-3.5 for Section 1.A. and GPT-4 for Section 1.C) to group their explanations into common themes, creating various categories. To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
@@ C.1. Micro-level Principles
In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 1.A.
We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases. This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment.
For Figure 7(b), we observe an opposite pattern in the LLM reasoning.
A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
Finally, Figure 7(c) contains the reasons for Figure 3 based on having a common hobby, a common color, or a common location. We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
@@
Figure S.7. Distribution of reasoning across categories for the networks of Figures 1(a), 2(a), and 3. The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies). The error bars correspond to 95% confidence intervals.